# Cross-positive linear maps, positive polynomials and sums of squares

### v1

@@ 4.3. Reverse Hölder inequality
Using $J_{1},J_{2},J_{3},J_{4}$ are as stated in Lemma 4.13 for $n=5$, we get $J_{5}=\frac{1}{27720}$, which is also in accordance with the formula in Lemma 4.13 for $n=5$.
@@
In all of the formulas above the difference from the case $n\geq 6$ is that integration intervals for $\phi_{4}$ and $\psi_{3}$ are $[0,2\pi]$ instead of $[0,\pi]$. However, since
@@
we can replace both intervals $[0,2\pi]$ with $[0,\pi]$. Since the integrands are precisely as in the case $n\geq 6$ with $n=5$, the formulas for $J_{7}$ from the case $n\geq 6$ hold also when applied to $n=5.$
Case 3: $n=4$. Note that the parametrizations of $x_{i}(\underline{\phi})$, $i=1,2,3$, and $y_{i}(\underline{\phi},\underline{\psi})$, $i=1,2$, are the same as in the case $n\geq 6$, while
@@
By (4.5), the integration intervals for $\phi_{3}$ and $\psi_{2}$ can be replaced by $[0,\pi]$ instead of $[0,2\pi]$. Since the integral is precisely as in the case $n\geq 6$ with $n=4$, the formula for $J_{4}$ from the case $n\geq 6$ holds also when applied to $n=4.$ The argument for $J_{5}$ following the same formula as in the case $n\geq 6$ for $n=4$ is the following computation:
00 0 $=\int_{T}({\tt z}_{1}+{\tt z}_{2}+{\tt z}_{3}+{\tt z}_{4})^{4}\;{\rm d}\sigma=\sum_{i=1}^{4}\int_{T}{\tt z}_{i}^{4}\;{\rm d}\sigma+4\sum_{i\neq j}\int_{T}{\tt z}_{i}^{3}{\tt z}_{j}\;{\rm d}\sigma+6\sum_{i<j}\int_{T}{\tt z}_{i}^{2}{\tt z}_{j}^{2}\;{\rm d}\sigma$
@@
Using $J_{1},J_{2},J_{3},J_{4}$ are as stated in Lemma 4.13 for $n=4$, we get $J_{5}=\frac{1}{9600}$, which is also as stated in Lemma 4.13 for $n=4$.
@@
By (4.5), the integration intervals for $\phi_{3}$ and $\psi_{2}$ can be replaced by $[0,\pi]$ instead of $[0,2\pi]$. The difference in the integrands in comparison to the case $n\geq 6$ is that every summand lacks the $\cos(\phi_{4})^{2}$ term, while some summands lack the $\cos(\psi_{3})^{2}$ term. However, looking at the computation for $n\geq 6$ both correspond to the term $B_{n-5,0}=\frac{1}{n-3}$. For $n=4$ this term is equal to 1, so the final formula for $J_{7}$ is the same as in the case $n\geq 6$ when applied to $n=4.$
Case 4: $n=3$. Note that the parametrizations of $x_{1}(\underline{\phi})$, $x_{2}(\underline{\phi})$ and $y_{1}(\underline{\phi},\underline{\psi})$ are the same as in the case $n\geq 6$, while
@@
where we included only terms with nonzero integrals in the computations above. By (4.5), the integration intervals for $\phi_{2}$ and $\psi_{1}$ can be replaced by $[0,\pi]$ instead of $[0,2\pi]$. The integrands of $I_{2}$, $J_{2}$, $J_{6}^{(1)}$, $J_{6}^{(2)}$ are the same as in the case $n\geq 6$ and hence also the corresponding formulas when applied to $n=3$. The difference in the integrands of $J_{3}$ and $J_{4}$ in comparison with the case $n\geq 6$ is that some summands lack at least one of the terms $\cos(\phi_{3})^{2}$ or $\cos(\psi_{2})^{2}$. Looking at the computation for $n\geq 6$ these terms correspond to the factor $B_{n-4,0}=\frac{1}{n-2}$. For $n=3$ this term is equal to 1. So the final formulas for $J_{3}$ and $J_{4}$ are the same as in the case $n\geq 6$ when applied to $n=3$. Hence, also the formula for $J_{6}$ is the same when applied to $n=3$ by the argument as in the case $n\geq 6$. ∎
@@
(4.8) $\begin{split}&\sum_{i}d_{i}^{4}-\frac{1}{n-1}\sum_{i\neq j}4d_{i}^{3}d_{j}+2\left(1+\frac{2}{(n-1)^{2}}\right)\sum_{i<j}d_{i}^{2}d_{j}^{2}\\ &\hskip 28.45274pt-\frac{4}{n-1}\left(1-\frac{2}{n-1}\right)\sum_{\begin{subarray}{c}i,j,k\\ \text{pairw.diff},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}+\frac{1}{(n-1)^{2}}\sum_{i<j<k<l}24d_{i}d_{j}d_{k}d_{l}=\frac{1}{I_{1}^{2}}.\end{split}$
@@
(4.9) $\begin{split}0&\leq 9-\left(\sum_{i}d_{i}^{4}-\frac{\sum_{i\neq j}4d_{i}^{3}d_{j}}{n-1}\right)J_{1}-6\sum_{i<j}d_{i}^{2}d_{j}^{2}J_{3}-12\sum_{\begin{subarray}{c}i,j,k\\ \text{pairw.diff},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}J_{4}\\ &\hskip 142.26378pt-\sum_{i<j<k<l}24d_{i}d_{j}d_{k}d_{l}\cdot\left\{\begin{array}[]{rr}J_{5}&\text{if }n\geq 4,\\[3.00003pt] 0,&\text{if }n=3.\end{array}\right.\\ &=9-\frac{J_{1}}{I_{1}^{2}}+\sum_{i<j}d_{i}^{2}d_{j}^{2}\Big{(}-6J_{3}+2J_{1}+\frac{4}{(n-1)^{2}}J_{1}\Big{)}\\ &\hskip 56.9055pt+\sum_{\begin{subarray}{c}i,j,k\\ \text{pairw.diff},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}\Big{(}-12J_{4}-\frac{4}{n-1}J_{1}+\frac{8}{(n-1)^{2}}J_{1}\Big{)}\\ &\hskip 56.9055pt+\sum_{i<j<k<l}24d_{i}d_{j}d_{k}d_{l}\cdot\left\{\begin{array}[]{rr}-J_{5}+\frac{1}{(n-1)^{2}}J_{1},&\text{if }n\geq 4,\\[3.00003pt] 0,&\text{if }n=3.\end{array}\right.\\ &=9-\frac{J_{1}}{I_{1}^{2}}+\frac{12}{(n-1)^{2}n(n+1)(n+4)(n+6)}\\ &\hskip 56.9055pt\Big{(}(n-3)\Big{(}\sum_{i<j}d_{i}^{2}d_{j}^{2}(n-2)-2\sum_{\begin{subarray}{c}i,j,k\\ \text{pairw.diff},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}\Big{)}+12\sum_{i<j<k<l}d_{i}d_{j}d_{k}d_{l}\Big{)}.\end{split}$
@@
(4.13) $(n-3)\Big{(}\sum_{i<j}d_{i}^{2}d_{j}^{2}(n-2)-2\sum_{\begin{subarray}{c}i,j,k\\ \text{pairw.diff},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}\Big{)}+12\sum_{i<j<k<l}d_{i}d_{j}d_{k}d_{l}\geq 0.$
@@
So we are proving
(4.18) $(n-2)\Big{(}\sum_{i<j\leq n}d_{i}^{2}d_{j}^{2}(n-1)-2\sum_{\begin{subarray}{c}i,j,k\leq n\\ \text{pairw.diff},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}\Big{)}+12\sum_{i<j<k<l\leq n}d_{i}d_{j}d_{k}d_{l}\geq 0.$
@@
(4.19) $\begin{split}&(n-3)\Big{(}\sum_{i<j\leq n}d_{i}^{2}d_{j}^{2}(n-2)-2\sum_{\begin{subarray}{c}i,j,k\leq n\\ \text{pairw.diff},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}\Big{)}+12\sum_{i<j<k<l\leq n}d_{i}d_{j}d_{k}d_{l}\\ &\hskip 56.9055pt+2\Big{(}\sum_{i<j\leq n}d_{i}^{2}d_{j}^{2}(n-2)-\sum_{\begin{subarray}{c}i,j,k\leq n\\ \text{pairw.diff},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}\Big{)}\geq 0.\end{split}$
The first summand in (4.19) is nonnegative by the induction hypothesis, while the second by Muirhead’s inequality [Mui03] used for tuples $(2,2,\underbrace{0,\ldots,0}_{n-2})\succ(2,1,1,\underbrace{0,\ldots,0}_{n-3})$.
The moreover part follows by taking $g({\tt z})=\frac{1}{\sqrt{I_{1}}}{\tt z}_{1}$ and noticing that
@@
As $n\to\infty$, we have $\|g\|_{4}\to\sqrt{3}\|g\|_{2}$. ∎
@@
By the action $g({\tt x},{\tt y})=g(U^{-1}{\tt x},U^{-1}{\tt y})$, $U\in SO(n)$, we can assume without loss of generality (see [HJ13, Corollary 2.5.11] that $g$ is of the form
@@
The moreover part follows by taking $g({\tt w})=\frac{1}{\sqrt{I_{5}}}{\tt w}_{12}$ and noticing that
@@
As $n\to\infty$, we have $\|g\|_{4}\to\sqrt[4]{6}\|g\|_{2}$. ∎
@@
But the values of both terms are equal to 00 by Lemma 4.13, since at least one of the indices $i,j,k$ occurs an odd number of times.
@@
If $(i,j)\neq(k,l)$, then in ${\tt z}_{ij}{\tt z}_{kl}$, ${\tt z}_{ij}{\tt z}_{lk}$, ${\tt z}_{ji}{\tt z}_{kl}$, ${\tt z}_{ji}{\tt z}_{lk}$ at least of the indices $i,j,k,l$ occurs an odd number of times and hence the corresponding integral is equal to 0 by Lemma 4.13. If $(i,j)=(k,l)$, then ${\tt z}_{ij}{\tt z}_{kl}={\tt z}_{ij}^{2}$, ${\tt z}_{ij}{\tt z}_{lk}={\tt z}_{ij}{\tt z}_{ji}$, ${\tt z}_{ji}{\tt z}_{kl}={\tt z}_{ji}{\tt z}_{ij}$, ${\tt z}_{ji}{\tt z}_{lk}={\tt z}_{ji}^{2}$, and hence
@@
This proves Claim.∎
@@
where in the first inequality we used the triangle inequality for $\|\cdot\|_{4}$, in the second the statement of Proposition 4.11 for symmetric (resp. skew-symmetric) bilinear forms, in the third $4]{6}<\sqrt{3}$ and in the last $\|g_{s}\|_{2}+\|g_{a}\|_{2}\leq\sqrt{2}\|g\|_{2}.$ Indeed, by Claim we have

### v2

@@ 4.3. Reverse Hölder inequality
Using $J_{1},J_{2},J_{3},J_{4}$ as stated in Lemma 4.12 for $n=5$, we get $J_{5}=\frac{1}{27720}$, which is also in accordance with the formula in Lemma 4.12 for $n=5$.
@@
In all of the formulas above the difference from the case $n\geq 6$ is that integration intervals for $\phi_{4}$ and $\psi_{3}$ are $0,2\pi]$ instead of $0,\pi]$. However, since
@@
we can replace both intervals $0,2\pi]$ with $0,\pi]$. Since the integrands are precisely as in the case $n\geq 6$ with $n=5$, the formulas for $J_{7}$ from the case $n\geq 6$ hold also when applied to $n=5.$
Case 3: $n=4$. Note that the parameterizations of $x_{i}(\underline{\phi})$, $i=1,2,3$, and $y_{i}(\underline{\phi},\underline{\psi})$, $i=1,2$, are the same as in the case $n\geq 6$, while
@@
By (4.8), the integration intervals for $\phi_{3}$ and $\psi_{2}$ can be replaced by $0,\pi]$ instead of $0,2\pi]$. Since the integral is precisely as in the case $n\geq 6$ with $n=4$, the formula for $J_{4}$ from the case $n\geq 6$ holds also when applied to $n=4.$ The argument for $J_{5}$ following the same formula as in the case $n\geq 6$ for $n=4$ is the following computation:
$0$ $=\int_{T}({\tt z}_{1}+{\tt z}_{2}+{\tt z}_{3}+{\tt z}_{4})^{4}\;{\rm d}\sigma=\sum_{i=1}^{4}\int_{T}{\tt z}_{i}^{4}\;{\rm d}\sigma+4\sum_{i\neq j}\int_{T}{\tt z}_{i}^{3}{\tt z}_{j}\;{\rm d}\sigma+6\sum_{i<j}\int_{T}{\tt z}_{i}^{2}{\tt z}_{j}^{2}\;{\rm d}\sigma$
@@
Using $J_{1},J_{2},J_{3},J_{4}$ as stated in Lemma 4.12 for $n=4$, we get $J_{5}=\frac{1}{9600}$, which is also as stated in Lemma 4.12 for $n=4$.
@@
By (4.8), the integration intervals for $\phi_{3}$ and $\psi_{2}$ can be replaced by $0,\pi]$ instead of $0,2\pi]$. The difference in the integrands in comparison to the case $n\geq 6$ is that every summand lacks the $\cos(\phi_{4})^{2}$ term, while some summands lack the $\cos(\psi_{3})^{2}$ term. However, looking at the computation for $n\geq 6$ both correspond to the term $B_{n-5,0}=\frac{1}{n-3}$. For $n=4$ this term is equal to 1, so the final formula for $J_{7}$ is the same as in the case $n\geq 6$ when applied to $n=4.$
Case 4: $n=3$. Note that the parameterizations of $x_{1}(\underline{\phi})$, $x_{2}(\underline{\phi})$ and $y_{1}(\underline{\phi},\underline{\psi})$ are the same as in the case $n\geq 6$, while
@@
where we included only terms with nonzero integrals in the computations above. By (4.8), the integration intervals for $\phi_{2}$ and $\psi_{1}$ can be replaced by $0,\pi]$ instead of $0,2\pi]$. The integrands of $I_{2}$, $J_{2}$, $J_{6}^{(1)}$, $J_{6}^{(2)}$ are the same as in the case $n\geq 6$ and hence also the corresponding formulas when applied to $n=3$. The difference in the integrands of $J_{3}$ and $J_{4}$ in comparison with the case $n\geq 6$ is that some summands lack at least one of the terms $\cos(\phi_{3})^{2}$ or $\cos(\psi_{2})^{2}$. Looking at the computation for $n\geq 6$ these terms correspond to the factor $B_{n-4,0}=\frac{1}{n-2}$. For $n=3$ this term is equal to 1. So the final formulas for $J_{3}$ and $J_{4}$ are the same as in the case $n\geq 6$ when applied to $n=3$. Hence, also the formula for $J_{6}$ is the same when applied to $n=3$ by the argument as in the case $n\geq 6$. ∎
@@
(4.11) $\begin{split}&\sum_{i}d_{i}^{4}-\frac{1}{n-1}\sum_{i\neq j}4d_{i}^{3}d_{j}+2\left(1+\frac{2}{(n-1)^{2}}\right)\sum_{i<j}d_{i}^{2}d_{j}^{2}\\ &\hskip 28.45274pt-\frac{4}{n-1}\left(1-\frac{2}{n-1}\right)\sum_{\begin{subarray}{c}i,j,k\\ \text{pairwise}\\ \text{different},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}+\frac{1}{(n-1)^{2}}\sum_{i<j<k<l}24d_{i}d_{j}d_{k}d_{l}=\frac{1}{I_{1}^{2}}.\end{split}$
@@
$0$ $\leq 9-\left(\sum_{i}d_{i}^{4}-\frac{\sum_{i\neq j}4d_{i}^{3}d_{j}}{n-1}\right)J_{1}-6\sum_{i<j}d_{i}^{2}d_{j}^{2}J_{3}-12\sum_{\begin{subarray}{c}i,j,k\\ \text{pairwise}\\ \text{different},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}J_{4}$
$\hskip 142.26378pt-\sum_{i<j<k<l}24d_{i}d_{j}d_{k}d_{l}\cdot\left\{\begin{array}[]{rr}J_{5}&\text{if }n\geq 4\\[3.00003pt] 0,&\text{if }n=3\end{array}\right$.
$=9-\frac{J_{1}}{I_{1}^{2}}+\sum_{i<j}d_{i}^{2}d_{j}^{2}\Big{(}-6J_{3}+2J_{1}+\frac{4}{(n-1)^{2}}J_{1}\Big{)}$
$\hskip 56.9055pt+\sum_{\begin{subarray}{c}i,j,k\\ \text{pairwise}\\ \text{different},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}\Big{(}-12J_{4}-\frac{4}{n-1}J_{1}+\frac{8}{(n-1)^{2}}J_{1}\Big{)}$
$\hskip 56.9055pt+\sum_{i<j<k<l}24d_{i}d_{j}d_{k}d_{l}\cdot\left\{\begin{array}[]{rr}-J_{5}+\frac{1}{(n-1)^{2}}J_{1},&\text{if }n\geq 4\\[3.00003pt] 0,&\text{if }n=3\end{array}\right$.
$=9-\frac{J_{1}}{I_{1}^{2}}+\frac{12}{(n-1)^{2}n(n+1)(n+4)(n+6)}$
$\hskip 56.9055pt\Big{(}(n-3)\Big{(}\sum_{i<j}d_{i}^{2}d_{j}^{2}(n-2)-2\sum_{\begin{subarray}{c}i,j,k\\ \text{pairwise}\\ \text{different},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}\Big{)}+12\sum_{i<j<k<l}d_{i}d_{j}d_{k}d_{l}\Big{)}.$
@@
(4.12) $(n-3)\Big{(}\sum_{i<j}d_{i}^{2}d_{j}^{2}(n-2)-2\sum_{\begin{subarray}{c}i,j,k\\ \text{pairwise}\\ \text{different},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}\Big{)}+12\sum_{i<j<k<l}d_{i}d_{j}d_{k}d_{l}\geq 0.$
@@
Using (4.13) in (4.12) it remains to prove
(4.14) $(n-2)\Big{(}\sum_{i<j\leq n}d_{i}^{2}d_{j}^{2}(n-1)-2\sum_{\begin{subarray}{c}i,j,k\leq n\\ \text{pairwise}\\ \text{different},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}\Big{)}+12\sum_{i<j<k<l\leq n}d_{i}d_{j}d_{k}d_{l}\geq 0.$
@@
(4.15) $\begin{split}&(n-3)\Big{(}\sum_{i<j\leq n}d_{i}^{2}d_{j}^{2}(n-2)-2\sum_{\begin{subarray}{c}i,j,k\leq n\\ \text{pairwise}\\ \text{different},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}\Big{)}+12\sum_{i<j<k<l\leq n}d_{i}d_{j}d_{k}d_{l}\\ &\hskip 56.9055pt+2\Big{(}\sum_{i<j\leq n}d_{i}^{2}d_{j}^{2}(n-2)-\sum_{\begin{subarray}{c}i,j,k\leq n\\ \text{pairwise}\\ \text{different},\\ j<k\end{subarray}}d_{i}^{2}d_{j}d_{k}\Big{)}\geq 0.\end{split}$
In (4.15), the first summand is nonnegative by the induction hypothesis, the second summand by (4.13) and the third summand by Muirhead’s inequality [27]. Namely, let $(s_{1},s_{2},s_{3},\ldots,s_{n})=(2,2,0,\ldots,0)$ and $(t_{1},t_{2},t_{3},t_{4},\ldots,t_{n})=(2,1,1,0,\ldots,0)$. Note $\sum_{i=1}^{n}s_{i}=\sum_{i=1}^{n}t_{i}$ and $\sum_{i=1}^{k}s_{i}\geq\sum_{i=1}^{k}t_{i}$ for $k=1,\ldots,n-1$. Since $d_{i}\geq 0$ for $i=1,\ldots,n$, [27] implies
(4.16) $\sum_{\sigma\in S_{n}}d_{1}^{s_{\sigma(1)}}d_{2}^{s_{\sigma(2)}}\cdots d_{n}^{s_{\sigma(n)}}\geq\sum_{\sigma\in S_{n}}d_{1}^{t_{\sigma(1)}}d_{2}^{t_{\sigma(2)}}\cdots d_{n}^{t_{\sigma(n)}},$
where $S_{n}$ stands for the symmetric group on $\{1,2,\ldots,n\}$. Note (4.16) is equivalent to
$\sum_{i<j\leq n}2(n-2)!\cdot d_{i}^{2}d_{j}^{2}\geq\sum_{\begin{subarray}{c}i,j,k\leq n\\ \text{pairwise}\\ \text{different},\\ j<k\end{subarray}}2(n-3)!\cdot d_{i}^{2}d_{j}d_{k},$
which implies the nonnegativity of the third summand in (4.15).
It remains to prove the moreover part of Proposition 4.10 for symmetric bilinear forms, i.e., the constant $\sqrt{3}$ in the inequality $\|g\|_{4}\leq\sqrt{3}\|g\|_{2}$ is asymptotically sharp as $n\to\infty$. For $g({\tt z})=\frac{1}{\sqrt{I_{1}}}{\tt z}_{1}$ note
@@
As $n\to\infty$, we deduce that $\|g\|_{4}\to\sqrt{3}\|g\|_{2}$. ∎
@@
By the action $g({\tt x},{\tt y})=g(U^{-1}{\tt x},U^{-1}{\tt y})$, $U\in SO(n)$, we can assume without loss of generality (see [21, Corollary 2.5.11] that $g$ is of the form
@@
It remains to prove the moreover part of Proposition 4.10 for skew-symmetric bilinear forms, i.e., the constant $\sqrt[4]{6}$ in the inequality $\|g\|_{4}\leq\sqrt[4]{6}\|g\|_{2}$ is asymptotically sharp as $n\to\infty$. For $g({\tt w})=\frac{1}{\sqrt{I_{5}}}{\tt w}_{12}$, note
@@
As $n\to\infty$, we deduce that $\|g\|_{4}\to\sqrt[4]{6}\|g\|_{2}$. ∎
@@
But the values of both terms are equal to 0 by Lemma 4.12, since at least one of the indices $i,j,k$ occurs an odd number of times.
@@
If $(i,j)\neq(k,l)$, then in ${\tt z}_{ij}{\tt z}_{kl}$, ${\tt z}_{ij}{\tt z}_{lk}$, ${\tt z}_{ji}{\tt z}_{kl}$, ${\tt z}_{ji}{\tt z}_{lk}$ at least one of the indices $i,j,k,l$ occurs an odd number of times and hence the corresponding integral is equal to 0 by Lemma 4.12. If $(i,j)=(k,l)$, then ${\tt z}_{ij}{\tt z}_{kl}={\tt z}_{ij}^{2}$, ${\tt z}_{ij}{\tt z}_{lk}={\tt z}_{ij}{\tt z}_{ji}$, ${\tt z}_{ji}{\tt z}_{kl}={\tt z}_{ji}{\tt z}_{ij}$, ${\tt z}_{ji}{\tt z}_{lk}={\tt z}_{ji}^{2}$, and hence
@@
This proves the Claim. $\Box$
@@
where in the first inequality we used the triangle inequality for $\|\cdot\|_{4}$, in the second the statement of Proposition 4.10 for symmetric (resp. skew-symmetric) bilinear forms, in the third $\sqrt[4]{6}<\sqrt{3}$ and in the last $\|g_{s}\|_{2}+\|g_{a}\|_{2}\leq\sqrt{2}\|g\|_{2}.$ Indeed, by the Claim